# On the complexity of matrix Putinar’s Positivstellensatz

### v1

@@ 3. Degree bounds for scalar polynomial inequalities
Case I. Suppose that $F^{(\xi)}(x)\leq\frac{3F_{\min}}{4}$. Then we know that $x\notin K$. Let $y\in K$ be the vector such that $\|x-y\|_{2}=d(x,K)$. Since $F(y)\succeq F_{\min}\cdot I_{\ell}$, we have that
$\begin{array}[]{ll}\frac{F_{\min}}{4}\leq F^{(\xi)}(y)-F^{(\xi)}(x)&\leq\max\limits_{s\in\bar{\Delta}}\|\nabla F^{(\xi)}(s)\|_{2}\|x-y\|_{2}\\ &\leq\frac{2d(2d-1)}{\sqrt{n}+1}\|x-y\|_{2}\max\limits_{s\in\bar{\Delta}}|F^{(\xi)}(s)|,\\ \end{array}$
where the last inequality is from Theorem 2.3. Note that $\max_{x\in\bar{\Delta}}|F^{(\xi)}|\leq\|F^{(\xi)}\|_{B,d}$ (cf. [13]). It follows from Lemma 3.2 that
$\frac{F_{\min}}{4}\leq\frac{2d(2d-1)}{\sqrt{n}+1}\|F^{(\xi)}\|_{B,d}\|x-y\|_{2}\leq 2d^{2}\|F\|_{B}\|x-y\|_{2},$
By the Lojasiewicz inequality (3.5), we have that
$G(x)\geq\kappa^{-1}\|x-y\|_{2}^{\eta}\geq\frac{1}{{8}^{\eta}\kappa d^{2\eta}}(\frac{F_{\min}}{\|F\|_{B}})^{\eta}=\delta.$
Hence, there exists $i_{0}\in\{1,\dots,m\}$ such that $g_{i_{0}}(x)\leq-\delta$. By item (ii), we have $h_{i_{0}}(x)\geq\frac{1}{2}$. Since $\max_{x\in\bar{\Delta}}|F^{(\xi)}(x)|\geq F_{\min}$, we know that for $x\in\bar{\Delta}$,
$F^{(\xi)}(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)$ $\geq F^{(\xi)}(x)+\lambda\frac{1}{2}\delta-2\lambda v(m-1)$
$\geq F^{(\xi)}(x)+\lambda\frac{\delta}{4}+\lambda\left(\frac{\delta}{4}-2v(m-1)\right)$
$\geq-\|F^{(\xi)}(x)\|_{B,d}+\frac{5}{4}\|F^{(\xi)}(x)\|_{B,d}$
$\geq\frac{1}{4}\|F^{(\xi)}(x)\|_{B,d}\geq\frac{1}{4}\max_{x\in\bar{\Delta}}|F^{(\xi)}(x)|$
$\geq\frac{1}{4}F_{\min}.$
Case II. Suppose that $F^{(\xi)}(x)\geq\frac{3F_{\min}}{4}$. Then for every $x\in\bar{\Delta}$, the following holds
@@
Thus, we know that $F^{(\xi)}(x)\geq\frac{1}{4}F_{\min}>0$ for all $x\in\bar{\Delta}$. Note that $\xi\in\mathbb{R}^{n}$ with $\|\xi\|_{2}=1$ is arbitrary. We have $P\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ on $\bar{\Delta}$.
Since $\deg(h_{i})\leq 10^{5}nd_{G}^{2}\delta^{-2}v^{-1}$ $(i=1,\dots,m)$, we have that
@@
By Lemma 3.2, there exists $\xi_{0}\in\mathbb{R}^{n}$ with $\|\xi_{0}\|_{2}=1$ such that $\|P\|_{B}=\|\xi_{0}^{T}P\xi_{0}\|_{B,\deg(P)}$. Then, we get that
$\begin{array}[]{ll}\|P\|_{B}&=\|\xi_{0}^{T}P\xi_{0}\|_{B,\deg(P)}\\ &=\|F^{(\xi_{0})}-\lambda\sum_{i=1}^{m}h_{i}g_{i}\|_{B,\deg(P)}\\ &\leq\|F^{(\xi_{0})}\|_{B,\deg(P)}+\lambda\sum_{i=1}^{m}\|h_{i}g_{i}\|_{B,\deg(P)}\\ &\leq\|F^{(\xi_{0})}\|_{B,\deg(F)}+\lambda\sum_{i=1}^{m}\|h_{i}\|_{B,\deg(P)-\deg(g_{i})}\|g_{i}\|_{B,\deg(g_{i})}\\ &\leq\|F\|_{B}+\lambda m\leq(1+5\delta^{-1}m)\|f\|_{B}\\ &\leq{8}^{\eta+1}m\kappa d^{2\eta}(\frac{\|F\|_{B}}{F_{\min}})^{\eta}\|F\|_{B}.\\ \end{array}$
It follows from Lemma 3.1 that for
@@
all coefficient matrices of $P(x)$ in the Bernstein basis $\{B_{\alpha}^{(k)}(x):\alpha\in\mathbb{N}_{k}^{n}\}$ are positive definite, i.e., there exist matrices $P_{\alpha}^{(k)}\succ 0$ such that
@@
Since $1+x_{1},\cdots,1+x_{n},\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]$ and $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$, the conclusion holds for sufficiently large $C$. ∎∎
Remark. If the linear independence constraint qualification holds at every $x\in K$, it was shown in [2, 3] that the Lojasiewicz exponent as in (3.5) is equal to 1, i.e., $\eta=1$. Then, under the assumptions of Theorem 3.3, we know that $f\in\mbox{QM}[G]_{k}^{\ell}$ if
4. A scalar representation for polymomial matrix inequalities
In this section, we provide exact estimates on the quantity and degree bound for polynomials $d_{1},\dots,d_{t}$ with $d_{1},\dots,d_{t}\in\mbox{QM}[G]$ such that the polynomial matrix inequality $G(x)\succeq 0$ can be equivalently described by scalar polynomial inequalities $d_{1}\geq 0$, $\dots$, $d_{t}\geq 0$. This is a quantitative version of the following proposition due to Schmüdgen [45].
Proposition 4.1 (Proposition 9, [45]).
Let $G(x)$ be an $m$-by-$m$ symmetric polynomial matrix and $G(x)\neq 0$. Then there exist $m$-by-$m$ diagonal matrices $D_{r}$, $m$-by-$m$ symmetric matrices $X_{\pm,r}$ and polynomials $z_{r}\in$ $\Sigma[x],r=1,\ldots,t$, such that:
@@ 4. A scalar representation for polymomial matrix inequalities
    $X_{+r}X_{-r}=X_{-r}X_{+r}=z_{r}I,D_{l}=X_{-r}GX_{-r}^{T},z_{r}^{2}G=X_{+r}D_{r}X_{+r}^{T}$,
@@
    For $x\in\mathbb{R}^{n},A(x)\geq 0$ if and only if $D_{r}(x)\geq 0$ for all $r=1,\ldots,t$.
Let
(4.1) $K=\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\}.$
A direct corollary of Proposition 4.1 is that there exist finite many polynomials in $\mbox{QM}[G]$ such that $K$ can be equivalently described by these polynomials. Denote
@@
Recall that
$d_{G}=\max\{\deg(G_{ij}),~{}1\leq i,j\leq m\}.$
In the following, we provide a quantitative version of Proposition 4.1.
@@
$K=\{d_{1}\geq 0,\dots,d_{\theta(m)}\geq 0\}.$
Proof.
We prove the result by applying the induction on the matrix size $m$. When $m=2$, the inequality $G(x)=[G_{ij}(x)]_{i,j=1,2}\succeq 0$ can be equivalently described as
@@
We can verify that the above polynomials all belong to $\mbox{QM}[G]_{3d_{G}}$. Note that $\theta(2)=6$. Hence, it holds for $m=2$.
Suppose it is true for $m=k$ and consider $m=k+1$. For $i\in\{1,\ldots,n\}$, denote by $P_{i}$ the permutation matrix which permutes the first row and the $i$-th row. For $11\leq i<j\leq k+$ 1, denote the row operation matrix $Q^{(ij)}$ as
$(Q^{(ij)})_{uv}=\left\{\begin{array}[]{ll}1,&\text{if~{}}(u,v)=(1,1),\dots,(k+1,k+1)\text{~{}and~{}}(i,j),\\ 0,&\text{otherwise}.\\ \end{array}\right.$
Let
@@
One can check that for $11\leq i\leq j\leq k+$ 1, the above matrices $T_{ij}$ are all invertible and it holds that
@@
$s_{ii}=G_{ii}~{}(i=1,\dots,k+1),$
$s_{ij}=G_{ii}+G_{jj}+2G_{ij}~{}(1\leq i<j\leq k+1).$
Denote matrices
@@
Since $T_{ij}$ are scalar matrices, we have
$\deg(s_{ij})\leq d_{G},~{}\deg(B^{(ij)})\leq 3d_{G}.$
@@
If $x\in K$, i.e., $G(x)\succeq 0$, it follows from (4.3) that $G^{(ij)}(x)\succeq 0$, i.e.,
@@
For the converse part, suppose that $x\in\mathbb{R}^{n}$ satisfies (4.6). If $s_{ij}=0$ for all $11\leq i\leq j\leq k+$ 1, then we have that $G_{ij}(x)=0$ for all $11\leq i,j\leq k+$ 1, and thus $G(x)=0$. If there exist $11\leq i_{0}\leq j_{0}\leq k+$ 1 such that $s_{i_{0}j_{0}}(x)\neq 0$, then the relation (4.4) gives that
@@
Since $G^{(i_{0}j_{0})}(x)\succeq 0$, we know that $G(x)\succeq 0$.
By the induction, each $B^{(ij)}$ is a $k$-by-$k$ symmetric polynomial matrix with $\deg(B^{(ij)})\leq 3^{k-1}d_{G}$ for $11\leq i\leq j\leq k+$ 1 and there exist $d_{1}^{(ij)},\dots,d_{\theta(k)}^{(ij)}\in QM[G]_{3^{k}d_{G}}$ such that
$\{x\in\mathbb{R}^{n}\mid B^{(ij)}(x)\succeq 0\}=\{x\in\mathbb{R}^{n}\mid d_{1}^{(ij)}(x)\geq 0,\dots,d_{\theta(k)}^{(ij)}\geq 0\}.$
Combining with the relation (4.5), we have that
@@
The total number of the above constraints is
$\frac{(k+1)(k+2)}{2}+\frac{(k+1)(k+2)}{2}\theta(k),$
which equals to $\theta(k+1)$. Hence, we complete the proof.
Remark. There exist alternative ways to obtain finitely many scalar polynomial inequalities to describe $K$. For instance, consider the characteristic polynomial $det[\lambda I_{m}-G(x)]$ of $G(x)$. It can be expressed as
@@
Since $G(x)$ is symmetric, all eigenvalues are real and $det[\lambda I_{m}-G(x)]$ has only real roots and we know that $G(x)\succeq 0$ if and only if $g_{i}(x)\geq 0$ $(i=1,\dots,m)$. Hence, if polynomials $g_{2}(x),\dots,g_{m}(x)\in\mbox{QM}[G]$, the quantity $\theta(m)$ and the degree $3^{m-1}d_{G}$ as in Theorem 4.2 can be improved to $m$ and $md_{G}$, respectively. However, $g_{2}(x),\dots,g_{m}(x)$ typically do not belong to $\mbox{QM}[G]$.
5. Degree bounds for matrix Positivstellensatze

### v2

@@ 3 Degree bounds for scalar polynomial inequalities
Case I. Suppose that $F^{(\xi)}(x)\leq\frac{3F_{\min}}{4}$. Then, we know that $x\notin K$. Let $x^{*}\in K$ be a vector such that $\|x-x^{*}\|_{2}=d(x,K)$. Since $F(x^{*})\succeq F_{\min}\cdot I_{\ell}$, we have that
$\begin{array}[]{ll}\frac{F_{\min}}{4}\leq F^{(\xi)}(x^{*})-F^{(\xi)}(x)&\leq\max\limits_{s\in\bar{\Delta}}\|\nabla F^{(\xi)}(s)\|_{2}\|x-x^{*}\|_{2}\\ &\leq\frac{2d(2d-1)}{\sqrt{n}+1}\|x-x^{*}\|_{2}\max\limits_{s\in\bar{\Delta}}|F^{(\xi)}(s)|,\\ \end{array}$
where the last inequality follows from Theorem 2.3. Since $\max\limits_{s\in\bar{\Delta}}|F^{(\xi)}(s)|\leq\|F^{(\xi)}\|_{B,d}$ (see (3.3)), it holds that
(3.9) $F_{\min}\leq\max_{s\in\bar{\Delta}}|F^{(\xi)}(s)|\leq\|F^{(\xi)}\|_{B,d}\leq\|F\|_{B},$
where the last inequality is due to Lemma 3.3. Then, we know that
$\frac{F_{\min}}{4}\leq\frac{2d(2d-1)}{\sqrt{n}+1}\|F^{(\xi)}\|_{B,d}\|x-x^{*}\|_{2}\leq 2d^{2}\|F\|_{B}\|x-x^{*}\|_{2}.$
By the Łojasiewicz inequality (3.5), we have that
$\mathcal{G}(x)\geq\kappa^{-1}d(x,K)^{\eta}\geq\kappa^{-1}\|x-x^{*}\|_{2}^{\eta}\geq\frac{1}{{8}^{\eta}\kappa d^{2\eta}}(\frac{F_{\min}}{\|F\|_{B}})^{\eta}=\delta.$
From the definition of the function $\mathcal{G}(x)$ as in (3.4), there exists $i_{0}\in\{1,\dots,m\}$ such that $g_{i_{0}}(x)\leq-\delta$. Since $h_{i_{0}}(x)$ is an SOS (see item (iv)) and $g_{i_{0}}(x)\leq-\delta$, we have $h_{i_{0}}(x)\geq\frac{1}{2}$ by item (ii). Note that
$-h_{i_{0}}(x)g_{i_{0}}(x)\geq\frac{1}{2}\delta,~{}-h_{i}(x)g_{i}(x)\geq-2v\cdot\max\limits_{s\in\bar{\Delta}}|g_{i}(s)|\geq-2v\cdot\|g_{i}\|_{B}=-2v~{}(i=1,\dots,m).$
In the above, we use the inequality (3.3). Then, we have that
$F^{(\xi)}(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)$ $\geq F^{(\xi)}(x)+\lambda\cdot\frac{1}{2}\delta-2\lambda v(m-1)$
$\geq F^{(\xi)}(x)+\lambda\frac{\delta}{4}+\lambda\left(\frac{\delta}{4}-2v(m-1)\right).$
One can see from the relation (3.9) that $F^{(\xi)}(x)\geq-\|F^{(\xi)}\|_{B,d}$ and $F_{\min}\leq\|F\|_{B}$, which implies that
$\frac{\delta}{4}-2v(m-1)\geq\frac{\delta}{4}-2vm\geq\frac{\delta}{4}-\frac{\delta F_{\min}}{10\|F\|_{B}}\geq\frac{\delta}{4}-\frac{\delta}{10}>0,$
$\lambda\frac{\delta}{4}=\frac{5}{4}\|F\|_{B}\geq\frac{5}{4}\|F^{(\xi)}\|_{B,d}.$
Then, it follows that
$F^{(\xi)}(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)$ $\geq F^{(\xi)}(x)+\lambda\frac{\delta}{4}+\lambda\left(\frac{\delta}{4}-2v(m-1)\right)$
$\geq-\|F^{(\xi)}\|_{B,d}+\frac{5}{4}\|F^{(\xi)}\|_{B,d}$
$\geq\frac{1}{4}\|F^{(\xi)}\|_{B,d}\geq\frac{1}{4}\max_{s\in\bar{\Delta}}|F^{(\xi)}(s)|$
$\geq\frac{1}{4}F_{\min},$
where the second inequality follows from the inequality (3.3).
Case II. Suppose that $F^{(\xi)}(x)\geq\frac{3F_{\min}}{4}$. Note that $-h_{i}(x)g_{i}(x)\geq-2v\cdot\|g_{i}\|_{B}=-2v$ for $i=1,\dots,m$, and
$2m\lambda v=2m\cdot\frac{\delta F_{\min}}{20m\|F\|_{B}}\cdot 5\|F\|_{B}/\delta=\frac{1}{2}F_{\min}.$
Then, we have that
@@
This concludes the proof of (3.8).
Since $\xi\in\mathbb{R}^{n}$ with $\|\xi\|_{2}=1$ is arbitrary, we have $P(x)\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ on $\bar{\Delta}$. Note that $x$ can be any fixed point in $\bar{\Delta}$, we know that $P(x)\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ for all $x\in\bar{\Delta}$. Since $\deg(h_{i})\leq 10^{5}nd_{G}^{2}\delta^{-2}v^{-1}$ for $i=1,\dots,m$, we have that
@@
By Lemma 3.3, there exists $\xi_{0}\in\mathbb{R}^{n}$ with $\|\xi_{0}\|_{2}=1$ such that $\|P\|_{B}=\|\xi_{0}^{T}P\xi_{0}\|_{B,\deg(P)}$. Then, we obtain that
$\begin{array}[]{ll}\|P\|_{B}=\|\xi_{0}^{T}P\xi_{0}\|_{B,\deg(P)}&=\|F^{(\xi_{0})}-\lambda\sum_{i=1}^{m}h_{i}g_{i}\|_{B,\deg(P)}\\ &\leq\|F^{(\xi_{0})}\|_{B,\deg(P)}+\lambda\sum_{i=1}^{m}\|h_{i}g_{i}\|_{B,\deg(P)}\\ &\leq\|F^{(\xi_{0})}\|_{B,\deg(F)}+\lambda\sum_{i=1}^{m}\|h_{i}\|_{B,\deg(P)-\deg(g_{i})}\|g_{i}\|_{B,\deg(g_{i})}\\ &\leq\|F\|_{B}+\lambda m\leq(1+5\delta^{-1}m)\|F\|_{B}\\ &\leq{8}^{\eta+1}m\kappa d^{2\eta}(\frac{\|F\|_{B}}{F_{\min}})^{\eta}\|F\|_{B},\\ \end{array}$
where the second inequality is due to the inequality (3.3). Note that
$\begin{array}[]{ll}\deg(P)=\deg(F-\lambda\sum_{i=1}^{m}h_{i}g_{i}I_{\ell})&=\max\{\deg(F),\deg(h_{i}g_{i})~{}(i=1,\dots,m)\}\\ &\leq\max\{d,2\cdot 10^{6}\cdot{8}^{3\eta}mnd_{G}^{3}\kappa^{3}d^{6\eta}(\frac{\|F\|_{B}}{F_{\min}})^{3\eta+1}\}\\ &=2\cdot 10^{6}\cdot{8}^{3\eta}mnd_{G}^{3}\kappa^{3}d^{6\eta}(\frac{\|F\|_{B}}{F_{\min}})^{3\eta+1}.\end{array}$
It follows from Lemma 3.1 that all coefficient matrices of $P(x)$ in the Bernstein basis $\{B_{\alpha}^{(k)}(x):\alpha\in\mathbb{N}_{k}^{n}\}$ are positive definite, for
@@
i.e., there exist matrices $P_{\alpha}^{(k)}\succ 0$ such that
@@
In the following, we show that $\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2}$. Consider the optimization
(3.10) $\left\{\begin{array}[]{rl}\min&\sqrt{n}-x_{1}-\cdots-x_{n}\\ \mathit{s.t.}&1-\|x\|_{2}^{2}\geq 0.\\ \end{array}\right.$
One can verify that the optimal value of (3.10) is 0. Since $\sqrt{n}-x_{1}-\cdots-x_{n}$ and $-(1-\|x\|^{2})$ are SOS convex and the feasible set has an interior point, it follows from [32, Theorem 3.3] that the lowest order SOS relaxation of (3.10) is exact and the optimal values of all SOS relaxations are achievable. Consequently, we have that $\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2}$. Similarly, we know that $1+x_{1},\cdots,1+x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2}$. Hence, we have
$\left(\sqrt{n}-x_{1}-\cdots-x_{n}\right)^{k-|\alpha|}\left(1+x_{1}\right)^{\alpha_{1}}\cdots\left(1+x_{n}\right)^{\alpha_{n}}\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2(k-|\alpha|)+2\alpha_{1}+\dots+2\alpha_{n}}=\mbox{QM}[1-\|x\|_{2}^{2}]_{2k}.$
Then, from the definition of $B_{\alpha}^{k}(x)$, we have $B_{\alpha}^{k}(x)\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2k}$. Since $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]_{k_{0}}^{\ell}$ for some $k_{0}>0$, it follows that $1-\|x\|_{2}^{2}\in\mbox{QM}[G]_{k_{0}}$, and thus $B_{\alpha}^{k}(x)\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2k+k_{0}}$. Then, the conclusion holds for sufficiently large $C$. Note that
$K=\left\{x\in\mathbb{R}^{n}\mid\frac{g_{1}(x)}{\left\|g_{1}\right\|_{B}}\geq 0,\dots,\frac{g_{m}(x)}{\left\|g_{m}\right\|_{B}}\geq 0\right\}.$
It follows from Theorem 2.4 that $\eta\leq(d_{G}+1)(3d_{G})^{n+m-2}$.
4 A scalar representation for polymomial matrix inequalities
In this section, we provide exact estimates on the quantity and degree bounds for polynomials $d_{1},\dots,d_{t}$ with $d_{1},\dots,d_{t}\in\mbox{QM}[G]$, such that the polynomial matrix inequality $G(x)\succeq 0$ can be equivalently described by the scalar polynomial inequalities $d_{1}\geq 0$, $\dots$, $d_{t}\geq 0$. This serves as a quantitative version of the following proposition due to Schmüdgen [50].
Proposition 1 (Proposition 9, [50]).
Let $G(x)$ be an $m$-by-$m$ symmetric polynomial matrix and $G(x)\neq 0$. Then there exist $m$-by-$m$ diagonal polynomial matrices $D_{r}$, $m$-by-$m$ symmetric polynomial matrices $X_{\pm r}$ and polynomials $z_{r}\in$ $\Sigma[x],r=1,\ldots,t$, such that:
@@ 4 A scalar representation for polymomial matrix inequalities
    $X_{+r}X_{-r}=X_{-r}X_{+r}=z_{r}I,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D_{r}}=X_{-r}GX_{-r}^{T},z_{r}^{2}G=X_{+r}D_{r}X_{+r}^{T}$,
@@
    For $x\in\mathbb{R}^{n}$, $G(x)\succeq 0$ if and only if $D_{r}(x)\succeq 0$ for all $r=1,\ldots,t$.
Let $K=\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\}.$ A direct corollary of Proposition 1 is that there exist finitely many polynomials in $\mbox{QM}[G]$ such that $K$ can be equivalently described by these polynomials. We also refer to [7, Proposition 5] for a generalization of Proposition 1, where the author considers the case that $K$ can be described by possibly infinitely many polynomial matrix inequalities. Denote
@@
Recall that $d_{G}=\max\{\deg(G_{ij}),~{}1\leq i,j\leq m\}.$ In the following, we provide a quantitative version of Proposition 1.
@@
$K=\{x\in\mathbb{R}^{n}\mid d_{1}(x)\geq 0,\dots,d_{\theta(m)}(x)\geq 0\}.$
Proof 4.2.
We prove the result by applying induction on the matrix size $m$. Suppose that $m=2$ and $G(x)=[G_{ij}(x)]_{i,j=1,2}$. In the following, we show that the inequality $G(x)=[G_{ij}(x)]_{i,j=1,2}\succeq 0$ can be equivalently described as
@@
If $G(x)\succeq 0$, we know that $G_{11}(x)\geq 0$, $G_{22}(x)\geq 0$, $G_{11}(x)\cdot G_{22}(x)-G_{12}(x)^{2}\geq 0$, which imply the above inequalities. Suppose $x\in\mathbb{R}^{n}$ satisfies these inequalities. If $G_{11}(x)\neq 0$ ($G_{22}(x)\neq 0$, respectively), it follows from the inequality $G_{11}(x)\cdot(G_{11}(x)\cdot G_{22}(x)-G_{12}(x)^{2})\geq 0$  ($G_{22}(x)\cdot(G_{11}(x)\cdot G_{22}(x)-G_{12}(x)^{2})\geq 0$, respectively) that $G_{11}(x)\cdot G_{22}(x)-G_{12}(x)^{2}\geq 0$. This implies that $G(x)\succeq 0$. Otherwise, if $G_{11}(x)=G_{22}(x)=0$, the last inequality reduces to $-2G_{12}(x)^{3}\geq 0$ and the inequality $G_{11}(x)+2G_{12}(x)+G_{22}(x)\geq 0$ reduces to $G_{12}(x)\geq 0$. Then, we have that $G_{12}=0$ and $G(x)=0$. Note that
$\begin{bmatrix}G_{11}&0\\ -G_{12}&G_{11}\\ \end{bmatrix}\cdot G\cdot\begin{bmatrix}G_{11}&-G_{12}\\ 0&G_{11}\\ \end{bmatrix}=\begin{bmatrix}G_{11}^{3}&0\\ 0&G_{11}(G_{11}G_{22}-G_{12}^{2})\\ \end{bmatrix},$
$\begin{bmatrix}G_{22}&-G_{12}\\ 0&G_{22}\\ \end{bmatrix}\cdot G\cdot\begin{bmatrix}G_{22}&0\\ -G_{12}&G_{22}\\ \end{bmatrix}=\begin{bmatrix}G_{22}(G_{11}G_{22}-G_{12}^{2})&0\\ 0&G_{22}^{3}\\ \end{bmatrix},$
$A^{T}\cdot G\cdot A=\begin{bmatrix}(G_{11}+2G_{12}+G_{22})^{3}&0\\ 0&(G_{11}+2G_{12}+G_{22})((G_{11}+2G_{12}+G_{22})G_{22}-(G_{12}+G_{22})^{2})\\ \end{bmatrix},$
where
$A=\begin{bmatrix}1&0\\ 1&1\\ \end{bmatrix}\cdot\begin{bmatrix}G_{11}+2G_{12}+G_{22}&-G_{22}-G_{12}\\ 0&G_{11}+2G_{12}+G_{22}\\ \end{bmatrix}.$
We know that all the above scalar inequalities involve polynomials that belong to $\mbox{QM}[G]_{3d_{G}}$. Note that $\theta(2)=6$. Hence, the result holds for $m=2$.
Suppose the result holds for $m=k$ and consider the case that $m=k+1$. In the following, we construct block diagonal polynomial matrices with a $1\times 1$ block and a $k\times k$ block, such that the inequality $G(x)\succeq 0$ can be equivalently described by the main diagonal blocks of these matrices. This reduction in matrix size allows us to apply induction. For $i\in\{1,\ldots,n\}$, let $P_{i}$ denote the permutation matrix that swaps the first row and the $i$-th row. For $11\leq i<j\leq k+$ 1, denote the row operation matrix $Q^{(ij)}$ as
[…]
Note that $Q^{(ij)}G(Q^{(ij)})^{T}$ is the matrix obtained by first adding the $j$-th row of $G$ to the $i$-th row and then adding the $j$-th column to the $i$-th column. Let
@@
Clearly, for $i<j$, $T_{ij}GT_{ij}^{\mathrm{T}}$ is the matrix obtained by swapping the first row with the $i$-th row of $Q^{(ij)}G(Q^{(ij)})^{T}$ and then swapping the first column with the $i$-th column. Hence, we know that for $11\leq i\leq j\leq k+$ 1, the matrices $T_{ij}$ are all invertible and it holds that
@@
$s_{ii}=G_{ii}~{}(i=1,\dots,k+1),~{}s_{ij}=G_{ii}+G_{jj}+2G_{ij}~{}(1\leq i<j\leq k+1).$
Denote the matrices
@@
Since $T_{ij}$ are scalar matrices, we have $\deg(s_{ij})\leq d_{G},~{}\deg(B^{(ij)})\leq 3d_{G}.$
@@
If $x\in K$, i.e., $G(x)\succeq 0$, it follows from (4.2) that $G^{(ij)}(x)\succeq 0$, which implies that
@@
For the converse part, suppose $x\in\mathbb{R}^{n}$ satisfies (4.5). If $s_{ij}=0$ for all $11\leq i\leq j\leq k+$ 1, then we have $G_{ij}(x)=0$ for all $11\leq i,j\leq k+$ 1, leading to $G(x)=0$. If there exist $11\leq i_{0}\leq j_{0}\leq k+$ 1 such that $s_{i_{0}j_{0}}(x)\neq 0$, the relation (4.3) gives that
@@
Since $G^{(i_{0}j_{0})}(x)\succeq 0$, we conclude that $G(x)\succeq 0$.
Note that each $B^{(ij)}$ is a $k$-by-$k$ symmetric polynomial matrix with $\deg(B^{(ij)})\leq 3d_{G}$ for $11\leq i\leq j\leq k+$ 1. By induction, there exist $d_{1}^{(ij)},\dots,d_{\theta(k)}^{(ij)}\in QM[G]_{3^{k}d_{G}}$ such that
$\{x\in\mathbb{R}^{n}\mid B^{(ij)}(x)\succeq 0\}=\{x\in\mathbb{R}^{n}\mid d_{1}^{(ij)}(x)\geq 0,\dots,d_{\theta(k)}^{(ij)}(x)\geq 0\}.$
Combining this with the relation (4.4), we have that
@@
The total number of the above constraints is $\frac{1}{2}(k+1)(k+2)+\frac{1}{2}(k+1)(k+2)\theta(k),$ which equals to $\theta(k+1)$. Hence, we have completed the proof.
Remark. There exist alternative ways to obtain finitely many scalar polynomial inequalities that describe $K$. For instance, consider the characteristic polynomial $det[\lambda I_{m}-G(x)]$ of $G(x)$. It can be expressed as
@@
Since $G(x)$ is symmetric, all eigenvalues are real, and $det[\lambda I_{m}-G(x)]$ has only real roots. We know that $G(x)\succeq 0$ if and only if $g_{i}(x)\geq 0$ for $i=1,\dots,m$. Note that $g_{1}(x)=\operatorname{trace}(G(x))$, which is always in $\mbox{QM}[G]$. Hence, if the polynomials ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g_{2}(x)},\dots,g_{m}(x)\in\mbox{QM}[G]$, the quantity $\theta(m)$ and the degree $3^{m-1}d_{G}$ as stated in Theorem 4.1 can be improved to $m$ and $md_{G}$, respectively. However, $g_{2}(x),\dots,g_{m}(x)$ typically do not belong to $\mbox{QM}[G]$.
5 Degree bounds for matrix Positivstellensätze